# DART[Uncaptioned image]: An AIGT Detector using AMR of Rephrased Text

### v1

@@ Appendix A Prompts
A.1 DART’s rephraser
We used the following prompt in the rephraser.
Please rewrite the following paragraph in {n} words: {paragraph}
A.2 AIGT datasets
We followed the prompts used in SeqXGPT (Wang et al., 2023) when generating the AIGT dataset. As Llama and Gemini inserted some prefix sentences in the result (e.g., “Here is the generation of …"), we used additional prompts to ensure the generated text’s consistency and quality.
@@ A.2 AIGT datasets
We used the following prompts for GPT-3.5-turbo and GPT-4o, for datasets except PubMedQA.
@@
For Llama3 and Gemini-1.5
We used the following prompts for Llama and Gemini-1.5, for datasets except PubMedQA.
@@ B.1 Environment
Hardware configuration
The experiments were conducted on a system with an AMD Ryzen Threadripper 3960X 24-Core Processor and four NVIDIA RTX A6000 GPUs. The four NVIDIA RTX A6000 GPUs are used for training existing detectors and executing AMR parsers. The semantic gap scoring module was run on a single core of the CPU.
LLM APIs
We used commercial APIs of LLMs to collect AIGTs and rephrased texts. GPT models are called with OpenAI’s official API. Llama 3 is called with a free API provided by Groq.com. Lastly, Gemini is called with OpenRouter’s API.
@@
We used Python 3.11.7 for implementing DART[Uncaptioned image]. Using scikit-learn library, we implemented DARTSVM and DARTDT with SVC and DecisionTreeClassifier. We mostly used the basic settings of those classes without doing a hyperparameter search. The only exception is the depth of the tree in DARTDT, and we set it as 5 based on our heuristic.
@@ B.2 Dataset statistics
Table 3 in page 3 shows the statistics of the collected dataset. We used four datasets, which belong to different domains: XsumNarayan et al. (2018), SQuADRajpurkar et al. (2018), RedditFan et al. (2018), and PubMedQAJin et al. (2019). Xsum is a dataset of news articles and summaries representing the general domain. SQuAD is an academic essay dataset using Wikipedia articles. Reddit is a dataset of human-written stories with writing prompts associated with creative writing. PubMedQA, consisting of question-answer pairs in the medical domain, is used to observe performance in a specific domain.
The statistics show that the average lengths of texts in each dataset are different. For example, Gemini-1.5 usually generates the most extended text on the PubMedQA dataset, while the model generates the shortest text on the Xsum and Reddit datasets. On average, it seems that the length of a given text is not a significant factor for discriminating origin.
@@ Appendix C Additional analysis
C.1 PCA result
Figure 4 and 5 in page 4 display PCA plots of DARTDT in the multi-candidate experiment. The figures show that each class makes several clusters. We interpret DART’s experimental results in the multi-candidate experiment by analyzing the PCA results. The distribution of feature vectors may affect the performance of SVM and DT classifiers. As SVM seeks a global decision boundary that maximizes margin, SVM may not find a clear decision boundary between multiple mini clusters. Meanwhile, DT can split such mini clusters based on multiple criteria. So, DT could achieve high performance in discriminating AIGTs from human-written texts. For example, we can easily discriminate humans from others and iteratively build different decision boundaries between smaller clusters. That is why DARTDT showed higher performance than DARTSVM.
C.2 Detailed result of efficiency
@@
(c) Llama 3
(d) Gemini-1.5 Flash

### v2

@@ Appendix A Prompts
A.1 AIGT datasets
In general, we followed the prompts used in SeqXGPT (Wang et al., 2023) when generating the AIGT dataset. We collected AIGTs by providing LLMs with the first 30 tokens of human-written texts and letting them generate the rest of the texts, except for the PubMedQA dataset. Besides, we asked LLMs to answer the questions in the PubMedQA dataset instead of providing the 30 tokens of text, borrowing the collecting method of Mitchell et al. (2023). We used different methods for PubMedQA because most of the texts in PubMedQA were shorter than 30 tokens. In addition, to avoid collecting AIGTs with irrelevant phrases (e.g., “Here is the generation of …”), we added a constraint clause in the prompts for Llama 3-70B and Gemini-1.5 Flash.
We understand that different datasets and different prompting methods may affect the performance of the detectors. Therefore, we conducted additional per-subset experiments to investigate whether those differences influenced the detecting performance. The findings are detailed in Appendix C.2.
@@ A.1 AIGT datasets
When collecting AIGTs with GPT-3.5-turbo and GPT-4o, we used the following prompts except for the PubMedQA dataset.
@@
For Llama 3-70B and Gemini-1.5-Flash
When collecting AIGTs with Llama 3-70B and Gemini-1.5-Flash, we used the following prompts except for the PubMedQA dataset.
@@
A.2 DART’s rephraser
When rephrasing a text into another rephrased version, we used the following prompt in the rephraser module.
Please rewrite the following paragraph in {n} words: {paragraph}
We used this prompt because we observed some semantic meanings of rephrased texts were largely changed without any prompting method in our pre-experiment. For example, some rephrased texts were much longer or shorter than the original texts, which was enough to distort the core message of the origins. As such distortion leads to unintended trivial semantic differences, we wanted to avoid such too-short or too-long texts. Thus, we restricted the word counts of rephrased texts by using prompts. Table 4 on page 4 shows the average number of words in the original and rephrased texts that we collected. It shows that the number of words slightly changed after rephrasing. We believe that such changes are minor to affect the performance of DART.
$T_{0}$ $T_{1}$ $T_{2}$
Human 267.95 258.47 270.38
GPT-3.5-T 107.48 89.85 83.08
GPT-4o 260.03 253.59 262.56
Llama3 152.33 133.94 127.69
Gemini-1.5 131.32 116.74 110.25
Table 4: Average number of words after rephrasing
Mac F1 Xsum SQuad Reddit PubMed
SeqXGPT* 63.0 75.1 57.0 58.2 61.7
DARTSVM 88.8 80.0 92.4 93.2 89.8
   DARTDT 98.6 99.0 98.4 98.6 98.4
Table 5: Performance of AIGT detectors across different subsets in a Multi-Candidate setting
@@ B.1 Environment
Hardware configuration:
The experiments were conducted on a system with an AMD Ryzen Threadripper 3960X 24-Core Processor and four NVIDIA RTX A6000 GPUs. The four NVIDIA RTX A6000 GPUs are used to train existing detectors and execute AMR parsers. The semantic gap scoring module was run on a single core of the CPU.
LLM APIs:
We used commercial APIs of LLMs to collect AIGTs and rephrased texts. GPT models are called with OpenAI’s official API. Llama 3-70B is called with a free API provided by groq.com. Lastly, Gemini-1.5-Flash is called with OpenRouter’s API.
@@
We used Python 3.11.7 for implementing DART[Uncaptioned image]. Using scikit-learn library, we implemented DARTSVM and DARTDT with SVC and DecisionTreeClassifier. We mostly used the basic settings of those classes without conducting a hyperparameter search. The only exception is the depth of the pruned tree in DARTDT, and we set it as 5 based on our heuristic.
@@ B.2 Dataset statistics
Table 7 in page 7 shows the statistics of the collected dataset. We used four datasets, which belong to different domains: Xsum Narayan et al. (2018), SQuAD Rajpurkar et al. (2018), Reddit Fan et al. (2018), and PubMedQA Jin et al. (2019). Xsum is a dataset of news articles and summaries. SQuAD is a question-answering dataset whose questions are based on Wikipedia articles. Reddit is a dataset of human-written stories with writing prompts. PubMedQA is a question-answering dataset on a specialized medical domain.
The statistics show that the average lengths of texts in each dataset are different. For example, Gemini-1.5 usually generates long texts on the PubMedQA dataset, while the model generates short texts on the Xsum and Reddit datasets. On average, it seems that the length of a given text is not a significant factor for discriminating origin.
@@ Appendix C Additional analysis
C.1 Precision, Recall
$p_{1}$ $p_{2}$ $r_{1}$ $r_{2}$
Human 0.619 0.582 0.600 0.561
GPT-3.5-T 0.645 0.605 0.631 0.595
GPT-4o 0.636 0.596 0.623 0.587
Llama3 0.648 0.610 0.631 0.594
Gemini-1.5 0.651 0.615 0.633 0.596
Table 6: Precision and Recall values for text comparisons between $T_{0}$, $T_{1}$ and $T_{0}$, $T_{2}$
As we discussed in Section 3, DART computes precision $p$ and recall $r$ scores with SEMA. Note that $p_{i}$ and $r_{i}$ refer to the semantic similarity between the original text $T_{0}$ and the $i$-th rephrased text $T_{i}$. DART assumes that the differences between those rephrased texts in terms of $p$ and $r$ values can be used to identify AIGTs. In this section, we provide evidence that supports the assumption by comparing the trend of $p$ and $r$ values.
Table 6 on page 6 illustrates the average of precision and recall values we collected. On average, the table shows that $p_{2}$ and $r_{2}$ are smaller than $p_{1}$ and $r_{1}$, respectively. This indicates that $T_{2}$ was semantically far from $T_{0}$ than $T_{1}$. So, as we apply rephraser more times on $T_{0}$, the semantics of rephrased text becomes farther from $T_{0}$.
Also, the result shows that $p$ and $r$ values are lower in human-written texts than AIGTs. For example, human-written text showed $p_{1}$ of 0.619, which is lower than AIGTs (ranging from 0.636 to 0.651). So, it is reasonable to use these values to distinguish between human-written texts and AIGTs.
C.2 Effect of prompt and domain changes
Since we used different prompting methods and datasets in generating AIGTs, we conducted the per-subset experiment to investigate whether those differences affected the performance of detectors. Specifically, we conducted multi-candidate experiments for each subset. For example, instead of using all data, we trained and tested models only with texts from PubMedQA.
Table 5 on page 5 shows the results of the per-subset experiment. Though the domains and prompting methods are different across those subsets, DARTDT achieved consistently high-performance scores by showing 98.6% macro F1. Also, DARTSVM (ranging from 80.0% to 93.2%) showed better consistency than SeqXGPT (ranging from 57.0% to 75.1%). This result indicates that DART[Uncaptioned image] models are robust on changes of domains or prompting methods compared to SeqXGPT.
C.3 Principal components of features
Figure 4 and 5 in page 4 display PCA plots of features used in DART. The figures show that each source makes several clusters. Here, we attempt to interpret DART’s experimental results by analyzing the PCA results. The distribution of feature vectors may affect the performance of SVM and DT classifiers. As SVM seeks a global decision boundary that maximizes margin, SVM may not find a clear decision boundary between multiple mini clusters. Meanwhile, DT can split such mini clusters based on multiple criteria. So, DT could achieve high performance in discriminating AIGTs from human-written texts. For example, we can easily discriminate humans from others and iteratively build different decision boundaries between smaller clusters. As a result, DARTDT can clearly discriminate sources and showed higher performance than DARTSVM.
C.4 Training efficiency on single-candidate setting
@@
(c) Llama 3-70b
(d) Gemini-1.5-Flash